Compton Meeting, June 2, 2011

From PolWiki
Jump to: navigation, search

Previous Meeting: Compton Meeting, May 23, 2011

Following Meeting:

Agenda

Laser

  1. Final dates for RadCon support
  2. Transfer function measurement
  3. Modifications to the LabView program

DAQ

  1. Update on summer upgrade plans

Photon detector

  1. Replay:
    1. Pedestals
    2. Analyzer status
  2. PbWO4:
    1. Detector modification update

Electron detector

  1. Analysis
    1. Update on analysis



Minutes

Attendance: Amrendra, Don, Manolis, Juan Carlos, Seth, Mark (chair), Wouter (minutes)

  • Introduction Seth, working as REU summer student on LabView controls for laser polarization measurement
  • Laser
    1. Mark will create atlis for radcon support to break vacuum
    2. LabView (Seth)
      • got used to LabView during first week
      • laser polarization measurement when light comes out of cavity (rotate polarizer to measure linear polarization)
      • added pause button, added delay to flipper (get rid of early signal spike)
      • all channels on same graph, working on error bars now
    3. Postpone transfer function measurement until after Qweak collab meeting (June 22--24), now start on June 27, Mark will double-check for conflicts with SAD schedule
    4. Measure gain/linewidth/etc precisely for NIM paper on laser cavity
  • DAQ
    1. Spare CPU board (6100) arrived at W&M, will bring to JLab, will get used in electron test stand
  • Photon detector
    1. PbWO4 modifications
      • Big bolt on stand might interfere with thermos, or would have to reposition PLT on thermos
      • Thermos measurements 12 inch (z) by 10 inch (width) by 9 inch (tall)
      • Great big styrofoam box, hopefully only for test
      • No room for lead under detector, lowering stand seems lot of work, room for 1 inch lead, need to find narrow lead, see picture.
      • Thermal model of thermos: total power of Compton signal, is there a gradient?
      • Hamlet wants to install fibers only once, Juan Carlos will look up fibers
      • Gamma beam profiler system: quote on XY translation stage (travel range of 6 inches), needs to be developed during summer, two scintillators with small PMT
      • Monte Carlo: modify geant3 to write events, then throw those into geant4 model of detector
    2. Replay
      • Mark collected info on pedestals, should be accurate enough for replay, Wouter requests table with run range where which pedestal is valid, and divide by number of samples
      • Kent is interested in also seeing the beam on, laser on/off pedestals, in addition to what is posted in elog 119
      • Analyzer status: BCMs are working now, combined BCMs in progress, same procedure for BPMs so should be straight-forward porting
  • Electron detector
    1. Analyzer: added event mode reading, not yet possible to calculate asymmetries
      • Only Amrendra working on this, Don created some macros
      • Weighting not done correctly (?), normalization by current
      • Electron detector meeting Thursday 4pm